OpenStudy (unklerhaukus):

Solve the differential equation: \[(1-x^2)\frac{\text d y}{\text dx}+xy=3x\]

OpenStudy (lgbasallote):

i see linear?

OpenStudy (unklerhaukus):

\[(1-x^2)\frac{\text d y}{\text dx}+xy=3x\]\[\frac{\text d y}{\text dx}+\frac{x}{1-x^2}y=\frac{3x}{1-x^2}\] \[R(x)=e^{\int \frac{x}{1-x^2}\text dx }=e^{\frac{-1}{2}\int \frac{-2x}{1-x^2}\text dx}=e^{\frac{-1}2\ln |1-x^2|}=\frac 1{\sqrt{1-x^2}}\] \[\frac{\text d }{\text dx}\left(\frac 1{\sqrt{1-x^2}}\cdot y\right)=\frac 1{\sqrt{1-x^2}}\cdot \frac{3x}{1-x^2}\]\[\frac{\text d }{\text dx}\left(\frac y{\sqrt{1-x^2}}\right)=\frac{3x}{(1-x^2)^{3/2}}\]\[\frac {y}{\sqrt{1-x^2}}=\int \frac{3x}{(1-x^2)^{3/2}}\text dx\]

OpenStudy (lgbasallote):

sounds right..

OpenStudy (lgbasallote):

\[\int \frac{3x}{(1-x^2)^{3/2}} dx\] that'sintegrable by u-sub right?

OpenStudy (unklerhaukus):

\[\frac {y}{\sqrt{1-x^2}}=\int \frac{3x}{(1-x^2)^{3/2}}\text dx\] \[\frac {y}{\sqrt{1-x^2}}=\frac{3}{\sqrt{1-x^2}}+c_1\]

OpenStudy (lalaly):

yes i got the same answer as u did

OpenStudy (lalaly):

\[y=3+c \sqrt{1-x^2}\]

OpenStudy (lgbasallote):

yup sounds right to me ^_^